Main INDEX
Monthly INDEX
PREV
NEXT
User name J. Singh
Log entry time 03:41:28 on August19,2003
Entry number 111873
keyword=status + owl shift summary , part 1
2330 CHL crash
0002 helium flow to hall (hall C as well) impeded? spectrometer magnets trip off. left dipole undergoes a fast dump. right dipole quenched while ramping down. septum was already off, so no major issues for it.
0030 Todd comes in for controlled access.
0100 Todd says that magnet reset went well. we can ramp the magnets back up when all the flows read above 60 l/m. Todd also talked to CHL about what the problem might be/have been. they said that they "lost a 1 of the 4 recovery compressors." apparently when this happens, it screws up the helium flow to the halls.
0145 informed by MCC that we're prob done for the night
0200 both upper and lower coils on septum are below 7K
0315 all flow levels for spec magnets are above 60 l/m. will wait for about 1 hr to insure that this flow is stable.
fig 1 is the current state of the magnets. the top strip is the flow in the left arm magnets. the bottom strip is the flow in the right arm magnets. these should all be above 60 l/m. the scale is from 0-100 l/m.
fig 2 is the septum stripchart since the beg of the shift.
orange = liquid helium reservoir level (60-90% range)
blue = upper coil temp (4-25K range)
red = lower coil temp (4-25K range)
green = yoke temp (4-25K range)
purple/pink = He gas flow (-20 l/m range)
FIGURE 1
FIGURE 2